# AdaRC: Mitigating Graph Structure Shifts during Test-Time

### v1

@@ 3.3 Adapting hop-aggregation parameters to restore representations
To mitigate the representation degradation caused by structure shifts, it becomes essential to adjust the featurizer of GNNs. In the following Proposition 3.5, we demonstrate that the degraded node representations due to structure shifts can be restored by adapting $\gamma$, the hop-aggregation parameter. This is because $\gamma$ determines the way to combine a node’s own attributes with its neighbors in GNNs. It is important to note that although our theory focuses on single-layer GCNs, a wide range of GNN models possess similar parameters for adaptation, e.g., the general PageRank parameters in GPRGNN [7], teleport probability in APPNP [17], layer aggregation in JKNet [43], etc.
Proposition 3.5 (Adapting $\gamma$).
Under the same learning setting as Proposition 3.4, adapting the source $\gamma_{S}$ to the optimal $\gamma_{T}=d_{T}(2h_{T}-1)$ on the target graph can alleviate the representation degradation and improve the target classification accuracy by $\Theta((\Delta h)^{2}+(\Delta d)^{2})$.
Proposition 3.5 indicates that the optimal $\gamma$ depends on both node degree $d$ and homophily $h$. For instance, consider a source graph with $h_{S}=1$ and $d_{S}=10$. In this case, the optimal featurizer assigns equal weight to the node itself and each of its neighbors, resulting in optimal $\gamma_{S}=10$. However, when the target graph’s degree remains unchanged but the homophily decreases to $h_{T}=0.5$, where each node’s neighbors are equally likely to be positive or negative, the neighbors no longer provide reliable information for node classification, leading to an optimal $\gamma_{T}=0$. Similarly, when the homophily remains the same, but the target graph’s degree is reduced to $d_{T}=1$, $\gamma_{S}$ overemphasizes the neighbor’s representation by placing excessive weight on it, whereas the optimal $\gamma_{T}$ in this case would be 1.
@@ 4 Proposed framework
  • How to ensure that our proposed algorithm is compatible with existing TTA algorithms in order to simultaneously address the co-existence of structure and attribute shifts?
In this section, we propose AdaRC, including a novel prediction-informed clustering loss to encourage high-quality node representations, and an adaptation framework compatible with a wide range of TTA algorithms. Figure 3 gives a general framework.
Figure 3: Our proposed framework of AdaRC (when combined with GPRGNN)
To adapt to graphs with different degree distributions and homophily, AdaRC uses GNNs that are capable of adaptively integrating multi-hop information, e.g., GPRGNN [7], APPNP [17], JKNet [43], etc. Specifically, we illustrate our framework using GPRGNN as a representative case. Notably, our framework’s applicability extends beyond this example, as demonstrated by the experimental results presented in Appendix C.8, showcasing its versatility across various network architectures.
@@ 4.1 Prediction-informed clustering loss
This subsection introduces how AdaRC updates the hop-aggregation parameters without labels. Previous TTA methods [22, 33, 46, 1] mainly adopt the entropy as a surrogate loss, as it measures the prediction uncertainty. However, we find that entropy minimization has limited effectiveness in improving representation quality (see Figure 4 and Table 6). Entropy is sensitive to the scale of logits rather than representation quality, often leading to trivial solutions. For instance, for a linear classifier, simply scaling up all the node representations can cause the entropy loss to approach zero, without improving the separability of the node representations between different classes. To address this issue, we propose the prediction-informed clustering (PIC) loss, which can better reflect the quality of node representation under structure shifts. Minimizing the PIC loss encourages the representations of nodes from different classes to be more distinct and less overlapping.
@@
It should be noted that although the form of PIC loss seems not reusing the adjacency matrix ${\bm{A}}$, it still evaluates the suitability of the current hop-aggregation parameters for the graph structure through the distribution of the representation ${\bm{Z}}$. As shown in Figure 4 and Proposition 3.4, structure shifts cause node representations to overlap more, leading to a smaller $\sigma^{2}_{\text{inter}}/\sigma^{2}_{\text{intra}}$ and a larger PIC loss. Alternatively, some algorithms, like SOGA [27], incorporate edge information by promoting connected nodes to share the same label. These designs implicitly assume of homophilic graph, limiting their applicability. As a result, SOGA performs poorly on heterophilic target graphs, as seen in Table 1. In contrast, our PIC loss directly targets GNN-encoded node representations, allowing it to generalize across different graph structures, whether homophilic or heterophilic.
By minimizing the PIC loss, we reduce intra-class variance while maximizing inter-class variance. Importantly, the ratio form of the PIC loss reduces sensitivity to the scale of representations; as the norm increases, the loss does not converge to zero, thus avoiding trivial solutions. It is also worth noting that the proposed PIC loss differs from the Fisher score [11] in two key aspects: First, PIC loss operates on model predictions, while Fisher score relies on true labels, making Fisher inapplicable in our setting where labels are unavailable. Second, PIC loss uses soft predictions for variance computation, which aids in the convergence of AdaRC, whereas the Fisher score uses hard labels, which can lead to poor convergence due to the unbounded Lipschitz constant, as we show in Theorem 4.1. We also provide an example in Appendix C.2 showing that AdaRC with PIC loss improves accuracy even when initial predictions are highly noisy.
@@ 4.2 Integration of generic TTA methods
Algorithm 1 AdaRC
0:   AdaRC (target graph ${\mathcal{T}}$, featurizer $f_{{\bm{\theta}},{\bm{\gamma}}}$, classifier $g_{{\bm{w}}}$)
@@
This subsection introduces how AdaRC integrates the adaptation of hop-aggregation parameters with existing TTA algorithms to simultaneously address the co-existence of structure and attribute shifts. Our approach is motivated by the complementary nature of adapting the hop-aggregation parameter and existing generic TTA methods. While the adapted hop-aggregation parameter effectively manages structure shifts, generic TTA methods handle attribute shifts in various ways. Consequently, we design a simple yet effective framework that seamlessly integrates the adaptation of hop-aggregation parameter with a broad range of existing generic TTA techniques.
Our proposed AdaRC framework is illustrated in Algorithm 1. Given a pre-trained source GNN model $f_{{\bm{\theta}},{\bm{\gamma}}}\circ g_{{\bm{w}}}$ and the target graph ${\mathcal{T}}$, we first employ the baseline TTA method, named BaseTTA, to produce the soft prediction $\hat{\bm{Y}}\in\mathbb{R}_{+}^{N\times C}$ as pseudo-classes, where $\sum_{c=1}^{C}\hat{\bm{Y}}_{i,c}=1$. Equipped with pseudo-classes, the hop-aggregation parameters ${\bm{\gamma}}$ is adapted by minimizing the PIC loss as described in Subsection 4.1. Intuitively, the predictions of BaseTTA are crucial for identifying pseudo-classes to cluster representations, and in return, better representations enhance the prediction accuracy of BaseTTA. Such synergy between representation quality and prediction accuracy mutually reinforces each other during the adaptation process, leading to much more effective outcomes. It is worth noting that AdaRC is a plug-and-play method that can seamlessly integrate with various TTA algorithms, including Tent [33], T3A [13], and AdaNPC [47].
@@
For each epoch, the computational complexity of the PIC loss is ${\mathcal{O}}(NCD)$, linear to the number of nodes. Compared to SOGA [27], which has quadratic complexity from comparing every node pair, PIC loss enjoys greater scalability to the graph size. For the whole AdaRC framework, it inevitably introduces additional computational overhead, which depends on both the GNN architecture and the baseline TTA algorithm. However, in practice, the additional computational cost is generally minimal since intermediate results (e.g. $\{{\bm{H}}^{(k)}\}_{k=0}^{K}$) can be cached and reused. We empirically evaluate the efficiency of AdaRC in Subsection 5.3.
@@
Finally, we analyze the convergence property of AdaRC in Theorem 4.1 below. The formal theorem and complete proofs can be found in Appendix B.
Theorem 4.1 (Convergence of AdaRC).
Let ${\bm{M}}=[\text{vec}({\bm{H}}^{(0)}),\cdots,\text{vec}({\bm{H}}^{(K)})]\in\mathbb{R}^{ND\times(K+1)}$ denote the concatenation of 0-hop to $K$-hop node representations. Given a base TTA algorithm, if (1) the prediction $\hat{{\bm{Y}}}$ is $L$-Lipschitz w.r.t. the (aggregated) node representation ${\bm{Z}}$, and (2) the loss function is $\beta$-smooth w.r.t. ${\bm{Z}}$, after $T$ steps of gradient descent with step size $\eta=\frac{1}{\beta\|{\bm{M}}\|_{2}^{2}}$, we have
@@
Theorem 4.1 shows that AdaRC is guaranteed to converge to a flat region with small gradients, with convergence rate $\frac{1}{T}$ and error rate $\propto L^{2}$. Essentially, the convergence of AdaRC depends on the sensitivity of the BaseTTA algorithm. Intuitively, if BaseTTA has large Lipschitz constant $L$, it is likely to make completely different predictions in each epoch, and thus hindering the convergence of AdaRC. However, in general cases, $L$ is upper bounded. We give theoretical verification in Lemma B.9 under ERM, and further empirically verify the convergence of AdaRC in Figure 6.
@@ 5 Experiments
We conduct extensive experiments on synthetic and real-world datasets to evaluate our proposed AdaRC from the following aspects:
  • RQ1: How can AdaRC empower TTA algorithms and handle various structure shifts on graphs?
  • RQ2: To what extent can AdaRC restore the representation quality better than other methods?
5.1 AdaRC handles various structure shifts (RQ1)
@@
We first adopt CSBM [8] to generate synthetic graphs with controlled structure and attribute shifts. We consider a hybrid of attribute shift, homophily shift and degree shift. For homophily shift, we generate a homophily graph with $h=0.8$ and a heterophily graph with $h=0.2$. For degree shift, we generate a high-degree graph with $d=10$ and a low-degree graph with $d=2$. For attribute shift, we transform the class centers ${\bm{\mu}}_{+},{\bm{\mu}}_{-}$ on the target graph. For real-world datasets, we adopt Syn-Cora [51], Syn-Products [51], Twitch-E [31], and OGB-Arxiv [12]. For Syn-Cora and Syn-Products, we use $h=0.8$ as the source graph and $h=0.2$ has the target graph. For Twitch-E and OGB-Arxiv, we delete a subset of homophilic edges in the target graph to inject both degree and homophily shifts. The detailed dataset statistics are provided in Appendix D.1.
We adopt GPRGNN [7] as the backbone model for the main experiments. We also provide results on other backbone models, including APPNP [17], JKNet [43], and GCNII [6] in Appendix C.8. Details on model architectures are provided in Appendix D.2. We run each experiment five times with different random seeds and report the mean accuracy and standard deviation.
@@
We consider two groups of base TTA methods, including: (1) generic TTA methods: T3A [13], Tent [33], and AdaNPC [47], and (2) graph TTA methods: GTrans [14], SOGA [27] and GraphPatcher [15]. To ensure a fair comparison, we focus on TTA algorithms in the same setting, which adapt a pre-trained model to a target graph without re-accessing the source graph. We adopt Empirical Risk Minimization (ERM) to pre-train the model on the source graph without adaptation. We use the node classification accuracy on the target graph to evaluate the model performance.
@@
The experimental results on the CSBM dataset are shown in Table 1. Under various shifts, the proposed AdaRC consistently enhances the performance of base TTA methods. Specifically, compared to directly using the pre-trained model without adaptation (ERM), adopting AdaRC (ERM+AdaRC) could significantly improve model performance, with up to 24.28% improvements. Compared with other baseline methods, AdaRC achieves the best performance in most cases, with up to 21.38% improvements. Besides, since AdaRC is compatible and complementary with the baseline TTA methods, we also compare the performance of baseline methods with and without AdaRC. As the results show, AdaRC could further boost the performance of TTA baselines by up to 22.72%.
For real-world datasets, the experimental results are shown in Table 2. Compared with ERM, AdaRC could significantly improve the model performance by up to 31.95%. Compared with other baseline methods, AdaRC achieves comparable performance on Twitch-E, and significant improvements on Syn-Cora, Syn-Products and OGB-Arxiv, with up to 40.61% outperformance. When integrated with other TTA methods, AdaRC can further enhance the performance by up to 39.31%. The significant outperformance verifies the effectiveness of the proposed AdaRC.
@@
In Appendix C.3 and C.4, we demonstrate that AdaRC exhibits robustness against (1) structure shifts of varying levels, and (2) additional adversarial shifts.
5.2 AdaRC restores the representation quality (RQ2)
Besides the superior performance of AdaRC, we are also interested in whether AdaRC successfully restores the quality of node representations under structure shifts. To explore this, we visualize the learned node representations on 2-class CSBM graphs in Figure 4. Although the pre-trained model generates high-quality node representations (Figure 4(a)), node representations degrades dramatically when directly deploying the source model to the target graph without adaptation (Figure 4(b)). With our proposed PIC loss, AdaRC successfully restores the representation quality with a clear cluster structure (Figure 4(f)). Moreover, compared to other common surrogate losses (entropy, pseudo-label), PIC loss results in significantly better representations.

### v2

@@ 3.3 Adapting hop-aggregation parameters to restore representations
To mitigate the representation degradation caused by structure shifts, it becomes essential to adjust the featurizer of GNNs. In the following Proposition 3.3, we demonstrate that the degraded node representations due to structure shifts can be restored by adapting $\gamma$, the hop-aggregation parameter. This is because $\gamma$ determines the way to combine a node’s own attributes with its neighbors in GNNs. Notice that although our theory mainly focuses on single-layer GCNs, a wide range of GNN models possess similar parameters for adaptation, e.g., the general PageRank parameters in GPRGNN (Chien et al., 2021), teleport probability in APPNP (Klicpera et al., 2019), layer aggregation in JKNet (Xu et al., 2018). We extend our analysis to multi-hop GCNs in Appendix A.7.
{restatable}
[Adapting $\gamma$]propositionpropadapt Under the same learning setting as Proposition 3.2, adapting the source $\gamma_{S}$ to the optimal $\gamma_{T}=d_{T}(2h_{T}-1)$ on the target graph can alleviate the representation degradation and improve the target classification accuracy by $\Theta((\Delta h)^{2}+(\Delta d)^{2})$.
Proposition 3.3 indicates that the optimal $\gamma$ depends on both node degree $d$ and homophily $h$. For instance, consider a source graph with $h_{S}=1$ and $d_{S}=10$. In this case, the optimal featurizer assigns equal weight to the node itself and each of its neighbors, resulting in optimal $\gamma_{S}=10$. However, when the target graph’s degree remains unchanged but the homophily decreases to $h_{T}=0.5$, where each node’s neighbors are equally likely to be positive or negative, the neighbors no longer provide reliable information for node classification, leading to an optimal $\gamma_{T}=0$. Similarly, when the homophily remains the same, but the target graph’s degree is reduced to $d_{T}=1$, $\gamma_{S}$ overemphasizes the neighbor’s representation by placing excessive weight on it, whereas the optimal $\gamma_{T}$ in this case would be 1. A visualization of these examples are given in Appendix A.2.
@@ 4 Proposed framework
  • How to ensure that our proposed algorithm is compatible with existing TTA algorithms (BaseTTA) in order to simultaneously address the co-existence of structure and attribute shifts?
In this section, we propose Matcha, including a novel prediction-informed clustering loss to encourage high-quality node representations, and an adaptation framework compatible with a wide range of TTA algorithms. Figure 3 gives a general framework.
Figure 3: Our proposed framework of Matcha (when combined with GPRGNN)
To adapt to graphs with different degree distributions and homophily, Matcha uses GNNs that are capable of adaptively integrating multi-hop information, e.g., GPRGNN (Chien et al., 2021), APPNP (Klicpera et al., 2019), JKNet (Xu et al., 2018), etc. Specifically, we illustrate our framework using GPRGNN as a representative case. Notably, our framework’s applicability extends beyond this example, as demonstrated by the experimental results presented in Appendix C.10, showcasing its versatility across various network architectures.
@@ 4.1 Prediction-informed clustering loss
This subsection introduces how Matcha updates the hop-aggregation parameters without labels. Previous TTA methods (Liang et al., 2020; Wang et al., 2021; Zhang et al., 2022; Bao et al., 2023) mainly adopt the entropy as a surrogate loss, as it measures the prediction uncertainty. However, we find that entropy minimization has limited effectiveness in improving representation quality (see Figure 4 and Table 7). Entropy is sensitive to the scale of logits rather than representation quality, often leading to trivial solutions. For instance, for a linear classifier, simply scaling up all the node representations can cause the entropy loss to approach zero, without improving the separability of the node representations between different classes. To address this issue, we propose the prediction-informed clustering (PIC) loss, which can better reflect the quality of node representation under structure shifts. Minimizing the PIC loss encourages the representations of nodes from different classes to be more distinct and less overlapping.
@@
It should be noted that although the form of PIC loss seems not reusing the adjacency matrix ${\bm{A}}$, it still evaluates the suitability of the current hop-aggregation parameters for the graph structure through the distribution of the representation ${\bm{Z}}$. As shown in Figure 4 and Proposition 3.2, structure shifts cause node representations to overlap more, leading to a smaller $\sigma^{2}_{\text{inter}}/\sigma^{2}_{\text{intra}}$ and a larger PIC loss. Alternatively, some algorithms, like SOGA (Mao et al., 2024), incorporate edge information by promoting connected nodes to share the same label. These designs implicitly assume of homophilic graph, limiting their applicability. As a result, SOGA performs poorly on heterophilic target graphs, as seen in Table 1. In contrast, our PIC loss directly targets GNN-encoded node representations, allowing it to generalize across different graph structures, whether homophilic or heterophilic.
By minimizing the PIC loss, we reduce intra-class variance while maximizing inter-class variance. Importantly, the ratio form of the PIC loss reduces sensitivity to the scale of representations; as the norm increases, the loss does not converge to zero, thus avoiding trivial solutions. It is also worth noting that the proposed PIC loss differs from the Fisher score (Gu et al., 2012) in two key aspects: First, PIC loss operates on model predictions, while Fisher score relies on true labels, making Fisher inapplicable in our setting where labels are unavailable. Second, PIC loss uses soft predictions for variance computation, which aids in the convergence of Matcha, whereas the Fisher score uses hard labels, which can lead to poor convergence due to the unbounded Lipschitz constant, as we show in Theorem 4.2. We also provide an example in Appendix C.2 showing that Matcha with PIC loss improves accuracy even when initial predictions are highly noisy.
@@ 4.2 Integration of generic TTA methods
Algorithm 1 Matcha
0:   Matcha (target graph ${\mathcal{T}}$, featurizer $f_{{\bm{\theta}},{\bm{\gamma}}}$, classifier $g_{{\bm{w}}}$, baseline TTA method BaseTTA)
@@
This subsection introduces how Matcha integrates the adaptation of hop-aggregation parameters with existing TTA algorithms to simultaneously address the co-existence of structure and attribute shifts. Our approach is motivated by the complementary nature of adapting the hop-aggregation parameter and existing generic TTA methods. While the adapted hop-aggregation parameter effectively manages structure shifts, generic TTA methods handle attribute shifts in various ways. Consequently, we design a simple yet effective framework that seamlessly integrates the adaptation of hop-aggregation parameter with a broad range of existing generic TTA techniques.
Our proposed Matcha framework is illustrated in Algorithm 1. Given a pre-trained source GNN model $f_{{\bm{\theta}},{\bm{\gamma}}}\circ g_{{\bm{w}}}$ and the target graph ${\mathcal{T}}$, we first employ the baseline TTA method, named BaseTTA, to produce the soft prediction $\hat{\bm{Y}}\in\mathbb{R}_{+}^{N\times C}$ as pseudo-classes, where $\sum_{c=1}^{C}\hat{\bm{Y}}_{i,c}=1$. Equipped with pseudo-classes, the hop-aggregation parameters ${\bm{\gamma}}$ is adapted by minimizing the PIC loss as described in Subsection 4.1. Intuitively, the predictions of BaseTTA are crucial for identifying pseudo-classes to cluster representations, and in return, better representations enhance the prediction accuracy of BaseTTA. Such synergy between representation quality and prediction accuracy mutually reinforces each other during the adaptation process, leading to much more effective outcomes. It is worth noting that Matcha is a plug-and-play method that can seamlessly integrate with various TTA algorithms, including Tent (Wang et al., 2021), T3A (Iwasawa & Matsuo, 2021), and AdaNPC (Zhang et al., 2023).
@@
For each epoch, the computational complexity of the PIC loss is ${\mathcal{O}}(NCD)$, linear to the number of nodes. Compared to SOGA (Mao et al., 2024), which has quadratic complexity from comparing every node pair, PIC loss enjoys greater scalability to the graph size. For the whole Matcha framework, it inevitably introduces additional computational overhead, which depends on both the GNN architecture and the baseline TTA algorithm. However, in practice, the additional computational cost is generally minimal since intermediate results (e.g. $\{{\bm{H}}^{(k)}\}_{k=0}^{K}$) can be cached and reused. We empirically evaluate the efficiency of Matcha in Subsection 5.3, Appendix C.8, and the scalability in Appendix C.9.
@@
Finally, we analyze the convergence property of Matcha in Theorem 4.2 below. The formal theorem and complete proofs can be found in Appendix B.
{restatable}
[Convergence of Matcha]theoremthmconverge Let ${\bm{M}}=[\text{vec}({\bm{H}}^{(0)}),\cdots,\text{vec}({\bm{H}}^{(K)})]\in\mathbb{R}^{ND\times(K+1)}$ denote the concatenation of 0-hop to $K$-hop node representations. Given a base TTA algorithm, if (1) the prediction $\hat{{\bm{Y}}}$ is $L$-Lipschitz w.r.t. the (aggregated) node representation ${\bm{Z}}$, and (2) the loss function is $\beta$-smooth w.r.t. ${\bm{Z}}$, after $T$ steps of gradient descent with step size $\eta=\frac{1}{\beta\|{\bm{M}}\|_{2}^{2}}$, we have
@@
Theorem 4.2 shows that Matcha is guaranteed to converge to a flat region with small gradients, with convergence rate $\frac{1}{T}$ and error rate $\propto L^{2}$. Essentially, the convergence of Matcha depends on the sensitivity of the BaseTTA algorithm. Intuitively, if BaseTTA has large Lipschitz constant $L$, it is likely to make completely different predictions in each epoch, and thus hindering the convergence of Matcha. However, in general cases, $L$ is upper bounded. We give theoretical verification in Lemma B.9 under ERM, and further empirically verify the convergence of Matcha in Figure 6.
@@ 5 Experiments
We conduct extensive experiments on synthetic and real-world datasets to evaluate our proposed Matcha from the following aspects:
  • RQ1: How can Matcha empower TTA algorithms and handle various structure shifts on graphs?
  • RQ2: To what extent can Matcha restore the representation quality better than other methods?
5.1 Matcha handles various structure shifts (RQ1)
@@
We first adopt CSBM (Deshpande et al., 2018) to generate synthetic graphs with controlled structure and attribute shifts. We consider a hybrid of attribute shift, homophily shift and degree shift. For homophily shift, we generate a homophily graph with $h=0.8$ and a heterophily graph with $h=0.2$. For degree shift, we generate a high-degree graph with $d=10$ and a low-degree graph with $d=2$. For attribute shift, we transform the class centers ${\bm{\mu}}_{+},{\bm{\mu}}_{-}$ on the target graph. For real-world datasets, we adopt Syn-Cora (Zhu et al., 2020), Syn-Products (Zhu et al., 2020), Twitch-E (Rozemberczki et al., 2021), and OGB-Arxiv (Hu et al., 2020). For Syn-Cora and Syn-Products, we use $h=0.8$ as the source graph and $h=0.2$ has the target graph. For Twitch-E and OGB-Arxiv, we delete a subset of homophilic edges in the target graph to inject both degree and homophily shifts. The detailed dataset statistics are provided in Appendix D.1.
We adopt GPRGNN (Chien et al., 2021) as the backbone model for the main experiments. We also provide results on other backbone models, including APPNP (Klicpera et al., 2019), JKNet (Xu et al., 2018), and GCNII (Chen et al., 2020) in Appendix C.10. Details on model architectures are provided in Appendix D.2. We run each experiment five times with different random seeds and report the mean accuracy and standard deviation.
@@
We consider two groups of base TTA methods, including: (1) generic TTA methods: T3A (Iwasawa & Matsuo, 2021), Tent (Wang et al., 2021), and AdaNPC (Zhang et al., 2023), and (2) graph TTA methods: GTrans (Jin et al., 2023), SOGA (Mao et al., 2024) and GraphPatcher (Ju et al., 2023). To ensure a fair comparison, we focus on TTA algorithms in the same setting, which adapt a pre-trained model to a target graph without re-accessing the source graph. We adopt Empirical Risk Minimization (ERM) to pre-train the model on the source graph without adaptation. We use the node classification accuracy on the target graph to evaluate the model performance.
@@
The experimental results on the CSBM dataset are shown in Table 1. Under various shifts, the proposed Matcha consistently enhances the performance of base TTA methods. Specifically, compared to directly using the pre-trained model without adaptation (ERM), adopting Matcha (ERM+Matcha) could significantly improve model performance, with up to 24.28% improvements. Compared with other baseline methods, Matcha achieves the best performance in most cases, with up to 21.38% improvements. Besides, since Matcha is compatible and complementary with the baseline TTA methods, we also compare the performance of baseline methods with and without Matcha. As the results show, Matcha could further boost the performance of TTA baselines by up to 22.72%.
For real-world datasets, the experimental results are shown in Table 2. Compared with ERM, Matcha could significantly improve the model performance by up to 31.95%. Compared with other baseline methods, Matcha achieves comparable performance on Twitch-E, and significant improvements on Syn-Cora, Syn-Products and OGB-Arxiv, with up to 40.61% outperformance. When integrated with other TTA methods, Matcha can further enhance the performance by up to 39.31%. The significant outperformance verifies the effectiveness of the proposed Matcha.
@@
We also demonstrate that Matcha exhibits robustness against (1) structure shifts of varying levels, (2) evolving target graph, and (3) additional adversarial shifts in Appendix C.3, C.4, and C.5, respectively.
5.2 Matcha restores the representation quality (RQ2)
Besides the superior performance of Matcha, we are also interested in whether Matcha successfully restores the quality of node representations under structure shifts. To explore this, we visualize the learned node representations on 2-class CSBM graphs in Figure 4. Although the pre-trained model generates high-quality node representations (Figure 4(a)), node representations degrades dramatically when directly deploying the source model to the target graph without adaptation (Figure 4(b)). With our proposed PIC loss, Matcha successfully restores the representation quality with a clear cluster structure (Figure 4(f)). Moreover, compared to other common surrogate losses (entropy, pseudo-label), PIC loss results in significantly better representations.